RSS Feed
Latest Updates
Jan
19
Software Updates
Posted by Jonathan Hughes on 19 January 2017 08:08 PM

We are currently performing some control panel software updates.

While this should not cause any service interruption there may be a brief period during any service restarts that are required.


Read more »



Jan
18
[RESOLVED] Loss of protection to DC08
Posted by Anthony Hogbin on 18 January 2017 05:47 PM

This afternoon we noticed an upstream fibre event.

Provider has confirmed a fibre break between Wrexham and St Asaph and is currently within SLA on predicted fix.

Networking team have confirmed a successful and seamless switch from one circuit to the other. There was no service interruption.

However until this circuit is repaired - the DC08 facility remains in a state of loss of protection.

We will update this ticket on works completion.

Thursday 19th January

0835hrs

The service has been returned, and protection has been returned to this location.


Read more »



Jan
18
[RESOLVED] Issues with Plesk55 / 217.194.212.23
Posted by Alex Gough on 18 January 2017 06:50 AM

We are currently aware of an issue affecting customers on our plesk55.myhostcp.com / 217.194.212.23 server at this time.

Engineers are currently working on resolving this issue at the moment.

Updates will be posted here.

**update**

Services have been restored, however engineers are still currently working on the server. Further updates to follow.

**Update**

The issue has now been marked as resolved. If you continue to have problems, please contact support with details of the issue.


Read more »